OpenStudy (anonymous):

DIFFERENTIATE Y=4SINX DIVIDE 2X+COSX

OpenStudy (anonymous):

@hartnn @mukushla @Miracrown

OpenStudy (anonymous):

you should know how to differentiate fractional functions like \(\frac{u}{v}\) ??

OpenStudy (anonymous):

use the formula: \[\left( \frac{u}{v} \right)'=\frac{u'v-v'u}{v^2}\]

OpenStudy (anonymous):

thank you